### v1

@@ Abstract
The advent of next-generation photometric and spectroscopic surveys is approaching, bringing more data with tighter error bars. As a result, theoretical models will become more complex, incorporating additional parameters, which will increase the dimensionality of the parameter space and make posteriors more challenging to explore. Consequently, the need to improve and speed up our current analysis pipelines will grow. In this work, we focus on the $3\times 2$ pt statistics, a summary statistic that has become increasingly popular in recent years due to its great constraining power. These statistics involve calculating angular two-point correlation functions for the auto- and cross-correlations between galaxy clustering and weak lensing. The corresponding model is determined by integrating the product of the power spectrum and two highly-oscillating Bessel functions over three dimensions, which makes the evaluation particularly challenging. Typically, this difficulty is circumvented by employing the so-called Limber approximation, which is an important source of error. We present Blast.jl, an innovative and efficient algorithm for calculating angular power spectra without employing the Limber approximation or assuming a scale-dependent growth rate, based on the use of Chebyshev polynomials. The algorithm is compared with the publicly available beyond-Limber codes, whose performances were recently tested by the LSST Dark Energy Science Collaboration in the N5K challenge. At similar accuracy, Blast.jl is $\approx 10$-15×15$\times$15 × faster than the winning method of the challenge, also showing excellent scaling with respect to various hyper-parameters.
@@ 1 Introduction
Next-generation spectroscopic and photometric surveys, such as Euclid (Mellier et al., 2024), LSST (Ivezic et al., 2019), the Nancy Grace Roman Space Telescope (Wang et al., 2022) and DESI (Levi et al., 2019), aim to achieve high-accuracy measurements of Large Scale Structure (LSS). These surveys measure various tracers, including galaxy density and weak lensing, across wider areas of the sky and at increasingly deep redshifts. Traditional cosmological parameter estimation relies on the power spectrum to compress the data into a reduced form, containing the majority of the desired information. For surveys that rely on photometric redshifts, there is little radial information, and the angular power spectra ($C_{\ell}$) within and between different tomographic redshift bins are sufficient to retain relevant information. This is the main idea behind $3\times 2$ pt statistics, which apply this technique for different combinations of tracers: galaxy clustering, lensing shear-shear correlation, and galaxy-galaxy lensing cross-correlation. A major drawback that comes with using $3\times 2$ pt statistics is the numerical computation of the model, which requires the calculation of a highly oscillating three-dimensional integral. A common work-around is the Limber approximation (Limber, 1953) which assumes that modes between structures at different epochs do not contribute to the line-of-sight integration. This approximation reduces the highly-oscillatory 3D integrand into a smooth 1D integrand, enabling faster computation. However, it introduces significant errors, particularly on large angular scales (low $\ell$’s). Therefore, a more accurate evaluation of the $3\times 2$ pt model is essential to better control systematics and to preserve the information gain from new surveys. We need not only a precise computation of the $3\times 2$ pt data vector but also a method that is as fast as possible. In fact, if this model is to be used to perform parameter inference through Monte Carlo Markov Chains (MCMC), we expect to perform millions of evaluations, so the task of speeding up the computation of the full non-Limber integral is very important as it will make cosmological analyses with $3\times 2$ pt computationally feasible. As an example, in the recent Euclid overview paper (Mellier et al., 2024), $3\times 2$ pt statistics were computed using the Limber approximation and the MCMC took $0.5$ million CPU hours to converge.
@@
In light of this problem, the LSST Dark Energy Science Collaboration (Ivezic et al., 2019) in 2020 launched the N5K challenge (Leonard et al., 2023), with the aim of assessing what the state of the art of non-Limber integration is, and determine which is the best method to incorporate in their pipeline. In this work we present a novel algorithm, implemented in the Blast.jl code, for non-Limber integration, as a late entry to the challenge. Blast.jl builds on the idea of performing a decomposition of the power spectrum on the convenient basis of the Chebyshev polynomials. Thanks to this decomposition, the hardest part of the integral becomes cosmology-independent and can therefore be pre-computed only once.
@@ 2.1 Angular Power Spectra
To derive the model for $3\times 2$ pt analysis, we begin by noting that any observable $A(\mathbf{x})$ defined on the sky can be expanded using spherical harmonics for scalar fields or spin-weighted spherical harmonics in the more general case. For scalar fields on the sphere, we restrict ourselves to the standard spherical harmonics, which form a complete and orthonormal basis for the total angular momentum operator, first introduced by Hu & White (1997). The expansion coefficients are:
$A_{\ell,m}=N(\ell)Y_{\ell,m}(\hat{n})e^{\mathrm{i}\mathbf{k\cdot x}},$ (1)
@@
where $N(\ell)$ is a $\ell$-dependent numerical prefactor and $Y_{\ell,m}(\hat{n})$ are the spherical harmonics. Then, if we have two scalar fields defined on the sphere $A(\mathbf{x})$ and $B(\mathbf{x})$, it is possible to expand both of them in spherical harmonics and compute the angular cross-correlation power spectrum:
$\begin{split}C_{\mathrm{AB}}(\ell)&=\langle A_{\ell,m}B_{\ell,m}^{*}\rangle=\\ &=N(\ell)\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})\frac{j_{\ell}(k\chi_{1})j_{\ell}(k\chi_{2})}{(k\chi_{1})^{\alpha}(k\chi_{2})^{\beta}}.\end{split}$ (2)
Here, $P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})$ is the $3$D power spectrum, containing the $2$ pt cosmological information, $\alpha$ and $\beta$ can assume different values depending on the probe that is considered, in particular 00 for clustering and $2$ for lensing, and $j_{\ell}(x)$ [^1] Relativistic effects such as Redshift Space Distortions and the Integrated Sachs-Wolfe effect introduce a dependency on the derivatives of the Bessel functions $j^{\prime}_{\ell}(x)$ and $j^{\prime\prime}_{\ell}(x)$. are the spherical Bessel functions, which are present in this expression due to the use of the plane wave expansion:
$e^{\mathrm{i}\mathbf{k\cdot r}}=4\pi\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}\mathrm{i}^{\ell}j_{\ell}(kr)Y_{\ell}^{m}(\hat{k})Y_{\ell}^{m*}(\hat{r}).$ (3)
Finally, we need to account for the fact that a real survey measures number counts and cosmic shear in different tomographic redshift bins. Those measurements are described by averaging over some survey-specific window functions $W_{i}^{A}(\chi)$ that only depend on geometry. The final expression for the angular power spectrum is then a triple integral of the form:
$\begin{split}C_{ij}^{\mathrm{AB}}(\ell)&=N(\ell)\int_{0}^{\infty}\mathrm{d}\chi_{1}W_{i}^{\mathrm{A}}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}W_{j}^{\mathrm{B}}(\chi_{2})\times\\ &\quad\times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})\frac{j_{\ell}(k\chi_{1})j_{\ell}(k\chi_{2})}{(k\chi_{1})^{\alpha}(k\chi_{2})^{\beta}}.\end{split}$ (4)
@@
Due to the nature of the spherical Bessel functions, which are highly oscillating and slowly damped, the evaluation of the integral is computationally challenging.
@@ 2.2 The Limber Approximation
$C^{\mathrm{AB}}_{ij}(\ell)\approx\int_{0}^{\infty}\frac{\mathrm{d}\chi}{\chi^{2}}K_{i}^{\mathrm{A}}(\chi)K_{j}^{\mathrm{B}}(\chi)P_{\mathrm{AB}}\left(k_{\ell},z\right).$ (7)
We defined $k_{\ell}\equiv(\ell+1/2)/f(\chi)$, $f(\chi)$ is a general function that depends on the cosmological model. In the case of the standard $\Lambda$CDM model, which we will assume here, $f(\chi)=\chi$. This integral is one dimensional and does not contain oscillatory function, making it inherently easier and faster to compute. The regimes of validity of this approximation are well understood (LoVerde & Afshordi, 2008). The Limber approximation, which sets $\chi_{1}=\chi_{2}$, works well especially for higher $\ell$’s, where the feature that the integrands have for $\chi_{1}=\chi_{2}$ gets increasingly sharp. In addition, for high multipoles $\ell$ the flat sky approximation is valid. Moreover, the approximation works best when the kernels $K_{i}^{\mathrm{A}}$ are broad in comoving distance (i.e., the approximation works best for the weak lensing sample rather than the clustering one) and for the auto-correlation within a redshift bin, as the kernels $K_{i}^{\mathrm{A}}$, $K_{j}^{\mathrm{A}}$ are superimposed in comoving distance.
@@
$\begin{split}C^{\mathrm{AB}}_{ij}(\ell)&=N(\ell)\int\frac{\mathrm{d}\chi}{\chi^{2}}K_{i}(\chi)K_{j}(\chi)P_{\mathrm{AB}}\left(k=\frac{\nu}{\chi}\right)\Biggl{\{}1+\\ &\quad+\frac{1}{\nu^{2}}\left[\frac{\chi^{2}}{2}\left(\frac{\tilde{K}_{i}^{{}^{\prime\prime}}(\chi)}{\tilde{K}_{i}(\chi)}+\frac{\tilde{K}_{j}^{{}^{\prime\prime}}(\chi)}{\tilde{K}_{j}(\chi)}\right)\right.+\\ &\quad\left.+\frac{\chi^{3}}{6}\left(\frac{\tilde{K}_{i}^{{}^{\prime\prime\prime}}(\chi)}{\tilde{K}_{i}(\chi)}+\frac{\tilde{K}_{j}^{{}^{\prime\prime\prime}}(\chi)}{\tilde{K}_{j}(\chi)}\right)\right]\Biggr{\}}.\end{split}$ (8)
@@ 3.1 The Algorithm
$\begin{split}C_{ij}^{gg}(\ell)&=\frac{2}{\pi}\int_{0}^{\infty}\mathrm{d}\chi_{1}W_{i}^{g}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}W_{j}^{g}(\chi_{2})\times\\ &\quad\times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})j_{\ell}(k\chi_{1})j_{\ell}(k\chi_{2}).\end{split}$ (10)
@@
$\begin{split}C_{ij}^{ss}(\ell)&=\frac{2(\ell+2)!}{\pi(\ell-2)!}\int_{0}^{\infty}\mathrm{d}\chi_{1}W_{i}^{s}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}W_{j}^{s}(\chi_{2})\times\\ &\quad\times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})\frac{j_{\ell}(k\chi_{1})}{(k\chi_{1})^{2}}\frac{j_{\ell}(k\chi_{2})}{(k\chi_{2})^{2}}.\end{split}$ (12)
@@
[…]
@@
$\begin{split}C_{ij}^{gs}(\ell)&=\frac{2}{\pi}\sqrt{\frac{(\ell+2)!}{(\ell-2)!}}\int_{0}^{\infty}\mathrm{d}\chi_{1}W_{i}^{g}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}W_{j}^{s}(\chi_{2})\times\\ &\quad\times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})\frac{j_{\ell}(k\chi_{1})}{(k\chi_{1})^{2}}j_{\ell}(k\chi_{2}).\end{split}$ (14)
@@
with
@@
However, to approximate $f(x)$, one can also use the polynomial obtained by interpolation in the Chebyshev points, a set of points defined as:
$x_{k}=\cos{\left(\frac{2k+1}{2n}\pi\right)}\,\,\,k=0,...,n-1.$ (19)
Those points are defined in $[-1,1]$ and have the property of being more dense close to the edges of the interval. The polynomial interpolated in $k$ Chebyshev points is:
$p_{k}(x)=\sum_{n=0}^{n_{\mathrm{max}}}c_{n}T_{n}(x).$ (20)
We highlight that $a_{n}$ and $c_{n}$ are different sets of coefficients (see the definitions in Eq. (17)-(20)). However, using the properties of Chebyshev polynomials, one can show (Trefethen, 2019) that $a_{n}$ and $c_{n}$ are related by:
$c_{n}=a_{n}+(a_{n+2k}+a_{n+4k}+...)+(a_{-n+2k}+a_{-n+4k}+...)$ (21)
for $11\leq n\leq k-$ 1. That is, the coefficients in the infinite series expansion are reassigned to their aliases of degree $<k$. This property is the reason why the Chebyshev polynomials are so widely used in approximation theory, as they are more accurate than any other method when fixing the number of interpolation points (Trefethen, 2019). In this specific problem, we approximate the power spectrum by interpolation in the Chebyshev points (Eq. (20)). This choice is driven by the fact that the coefficients $c_{n}(\chi_{1},\chi_{2};\theta)$ are evaluated through a Discrete Fourier Transform (DFT) (Press, 2007; Frigo & Johnson, 1997):
$c_{n}(\chi_{1},\chi_{2})=\sum_{j=0}^{N-1}P(k_{j},\chi_{1},\chi_{2})e^{2\mathrm{i}\pi jn/N}.$ (22)
The choice of the sampling points $k_{j}$ for the power spectrum will be motivated in the next section. Computing a FFT is much more efficient than evaluating the integrals defining the coefficients $a_{n}$ (Eq. (18)), which justifies our approach. The power spectrum, a (Lipschitz) continuous function over the interval $[k_{\text{min}},k_{\text{max}}]$ (which can be remapped in $[-1,1]$), in the basis of the Chebyshev polynomials becomes:
$P(k,\chi_{1},\chi_{2};\theta)\approx\sum_{n=0}^{n_{\text{max}}}c_{n}(\chi_{1},\chi_{2};\theta)T_{n}(k),$ (23)
where $\theta$ represents a generic set of cosmological parameters. This decomposition enables a clear separation of geometric and cosmological components in the integrals, simplifying the overall calculation. We can write:
@@
$f^{\mathrm{AB}}(k)=\begin{cases}k^{2}&\mathrm{AB}=gg\,,\\ 1/k^{2}&\mathrm{AB}=ss\,,\\ 1&\mathrm{AB}=gs\,.\end{cases}$ (26)
@@
$\begin{split}C_{ij}^{\mathrm{AB}}(\ell)&=\int_{0}^{\infty}d\chi\int_{0}^{1}dR\,\chi\left[\mathcal{K}_{i}^{\mathrm{A}}(\chi)\mathcal{K}_{j}^{\mathrm{B}}(R\chi)+\right.\\ &\quad\quad\left.+\mathcal{K}_{j}^{\mathrm{B}}(\chi)\mathcal{K}_{i}^{\mathrm{A}}(R\chi)\right]w_{\ell}^{\mathrm{AB}}(\chi,R\chi)\,,\end{split}$ (27)

### v2

@@ Abstract
[…]
@@ 1 Introduction
Next-generation spectroscopic and photometric surveys, such as Euclid (Mellier et al., 2024), Rubin Observatory Legacy Survey of Space and Time (LSST) (Ivezic et al., 2019), the Nancy Grace Roman Space Telescope (Wang et al., 2022) and Dark Energy Spectroscopic Instrument (DESI) (Levi et al., 2019), aim to achieve high-accuracy measurements of Large Scale Structure (LSS). These surveys measure various tracers, including galaxy density and weak lensing, across wider areas of the sky and at increasingly deep redshifts. Traditional cosmological parameter estimation relies on the power spectrum to compress the data into a reduced form, containing the majority of the desired information. For surveys that rely on photometric redshifts, there is little radial information, and the angular power spectra ($C_{\ell}$) within and between different tomographic redshift bins are sufficient to retain relevant information. This is the main idea behind $3\times 2$ pt statistics, which apply this technique for different combinations of tracers: galaxy clustering, lensing shear-shear correlation, and galaxy-galaxy lensing cross-correlation. A major drawback that comes with using $3\times 2$ pt statistics is the numerical computation of the model, which requires the calculation of a highly oscillating three-dimensional integral. A common work-around is the Limber approximation (Limber, 1953) which assumes that modes between structures at different epochs do not contribute to the line-of-sight integration. This approximation reduces the highly-oscillatory $3D$ integrand into a smooth $1D$ integrand, enabling faster computation. However, it introduces significant errors, particularly on large angular scales (low $\ell$’s). Therefore, a more accurate evaluation of the $3\times 2$ pt model is essential to better control systematics and to preserve the information gain from new surveys. We need not only a precise computation of the $3\times 2$ pt data vector but also a method that is as fast as possible. In fact, if this model is to be used to perform parameter inference through Monte Carlo Markov Chains (MCMC), we expect to perform millions of evaluations, so the task of speeding up the computation of the full non-Limber integral is very important as it will make cosmological analyses with $3\times 2$ pt computationally feasible. As an example, in the recent Euclid overview paper (Mellier et al., 2024), $3\times 2$ pt statistics were computed using the Limber approximation and the MCMC took $0.5$ million CPU hours to converge.
@@
In light of this problem, the LSST Dark Energy Science Collaboration (Ivezic et al., 2019) in 2020 launched the Non-local No-Nonsense Non-Limber Numerical Knockout (N5K) Challenge (Leonard et al., 2023), with the aim of assessing what the state of the art of non-Limber integration is, and determine which is the best method to incorporate in their pipeline. In this work we present a novel algorithm, implemented in the Blast.jl [^1] https://github.com/sofiachiarenza/Blast.jl code, for non-Limber integration, as a late entry to the challenge. Blast.jl builds on the idea of performing a decomposition of the power spectrum on the convenient basis of the Chebyshev polynomials. Thanks to this decomposition, the hardest part of the integral becomes cosmology-independent and can therefore be pre-computed only once.
@@ 2.1 Angular Power Spectra
To derive the model for the $3\times 2$ pt analysis, we start by noting that any observable can be expanded using spherical harmonics for scalar fields, such as the matter overdensity field $\delta(\mathbf{x})$. For more general cases, such as the lensing shear—a spin-$2$ field—spin-weighted spherical harmonics are used instead. In this section, we demonstrate how the angular power spectrum coefficients are derived for scalar fields on the sphere. The same approach applies to spin-$s$ fields, with the standard spherical harmonics $Y_{\ell,m}$ replaced by the spin-weighted spherical harmonics ${}_{s}Y_{\ell,m}$ (Zaldarriaga & Seljak, 1997; Chisari et al., 2019). The spherical harmonics, which provide a complete and orthonormal basis for the total angular momentum operator, were first introduced in this context by Hu & White (1997). A generic scalar quantity $\tilde{A}(\mathbf{\hat{n}})$, $\mathbf{\hat{n}}$ representing the unit vector pointing toward a given direction on the sky, is related to a $3$D quantity $A(\mathbf{r},z)$ through a projection kernel $W_{i}(z)$:
$\tilde{A}(\mathbf{\hat{n}})=\int\mathrm{d}zW_{i}(z)A(\chi(z)\mathbf{\hat{n}};z),$ (1)
$\chi(z)$ being the comoving distance. In terms of its Fourier transform, $A(\mathbf{r},z)$ can be expressed as:
$A(\chi(z)\mathbf{\hat{n}};z)=\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}A(\mathbf{k},z)e^{\mathrm{i}\mathbf{k\cdot\hat{n}}\,\chi(z)}.$ (2)
A plane wave can be expanded as:
$e^{\mathrm{i}\mathbf{k\cdot\hat{n}}r}=4\pi\sum_{\ell,m}\mathrm{i}^{\ell}j_{\ell}(kr)Y^{*}_{\ell,m}(\mathbf{\hat{k}})Y_{\ell,m}(\mathbf{\hat{n}}),$ (3)
this is also known as the Rayleigh’s formula. Inserting this expansion in Eq. (2), one obtains:
$A(\chi(z)\,\mathbf{\hat{n}};z)=4\pi\sum_{\ell,m}\mathrm{i}^{\ell}\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}A(\mathbf{k},z)\,j_{\ell}(k\chi(z))\\ \times Y^{*}_{\ell,m}(\mathbf{\hat{k}})Y_{\ell,m}(\mathbf{\hat{n}}).$ (4)
Putting the previous equations together, a generic quantity projected on the sky becomes:
$\tilde{A}(\mathbf{\hat{n}})=4\pi\sum_{\ell,m}\mathrm{i}^{\ell}\int\mathrm{d}z\,W_{i}(z)\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}A(\mathbf{k},z)\\ \times j_{\ell}(k\chi(z))Y^{*}_{\ell,m}(\mathbf{\hat{k}})Y_{\ell,m}(\mathbf{\hat{n}}).$ (5)
So, the coefficients of the spherical harmonics decomposition of $\tilde{A}(\mathbf{\hat{n}})$ are:
$A_{\ell,m}=4\pi\mathrm{i}^{\ell}\int\mathrm{d}z\,W_{i}(z)\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3}}A(\mathbf{k},z)\,j_{\ell}(k\chi(z))\\ \times Y^{*}_{\ell,m}(\mathbf{\hat{k}}).$ (6)
@@
Therefore, if we have two scalar fields defined on the sphere $\tilde{A}(\mathbf{n})$ and $\tilde{B}(\mathbf{n})$, it is possible to expand both of them in spherical harmonics and compute the angular cross-correlation power spectrum (Jeong et al., 2009; LoVerde & Afshordi, 2008; Hu & White, 1997):
$C_{\mathrm{AB}}(\ell)=\langle A_{\ell,m}B_{\ell,m}^{*}\rangle=\frac{2}{\pi}\int\mathrm{d}z\,W_{i}(z)\int\mathrm{d}z^{\prime}\,W_{j}(z^{\prime})\\ \times\int\mathrm{d}k\,k^{2}P_{AB}(k,z,z^{\prime})\,j_{\ell}(k\chi(z))j_{\ell}(k\chi(z^{\prime})).$ (7)
Here, $j_{\ell}(x)$ are the spherical Bessel functions [^2] Relativistic effects such as Redshift Space Distortions and the Integrated Sachs-Wolfe effect introduce a dependency on the derivatives of the Bessel functions $j^{\prime}_{\ell}(x)$ and $j^{\prime\prime}_{\ell}(x)$., which are present in this expression due to the use of the plane wave expansion, $P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})$ is the $3$D power spectrum, containing the $2$-point cosmological information, and the orthonormality condition of spherical harmonics is used. The power spectrum, as described in detail by Chisari et al. (2019), varies depending on the probes, since the transfer functions for matter overdensities and correlated galaxy shapes differ. However, these differences are very small, and, in practice, the matter power spectrum is commonly used. Following this convention, we adopt the matter power spectrum in this paper, consistent with the approach taken in the N5K challenge. $W_{i,j}^{AB}(z)$ are called window functions: they are survey-specific objects that only depend on the background cosmology and account for the fact that a real survey measures number counts and cosmic shear in different tomographic redshift bins.
Finally, accounting for probe-specific dependencies on $k$, $z$ and $\ell$, and changing variables from $z$ to $\chi(z)$ (the comoving distance), the angular power spectrum becomes:
$C_{ij}^{\mathrm{AB}}(\ell)=N(\ell)\int_{0}^{\infty}\mathrm{d}\chi_{1}\,W_{i}^{\mathrm{A}}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}\,W_{j}^{\mathrm{B}}(\chi_{2})\\ \times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})\frac{j_{\ell}(k\chi_{1})j_{\ell}(k\chi_{2})}{(k\chi_{1})^{\alpha}(k\chi_{2})^{\beta}}.$ (8)
@@
The exponents $\alpha$ and $\beta$ assume different values depending on the probe that is considered, in particular 00 for clustering and $2$ for lensing. Due to the nature of the spherical Bessel functions, which are highly oscillating and slowly damped, the evaluation of the integral is computationally challenging.
@@ 2.2 The Limber Approximation
$C^{\mathrm{AB}}_{ij}(\ell)\approx\int_{0}^{\infty}\frac{\mathrm{d}\chi}{\chi^{2}}K_{i}^{\mathrm{A}}(\chi)K_{j}^{\mathrm{B}}(\chi)P_{\mathrm{AB}}\left(k_{\ell},z\right),$ (11)
with
$k_{\ell}\equiv\frac{\ell+1/2}{\chi}.$ (12)
This formula is only valid in spatially flat cosmologies, which we assume throughout this paper. This integral is one dimensional and does not contain oscillatory function, making it inherently easier and faster to compute. The regimes of validity of this approximation are well understood (LoVerde & Afshordi, 2008). The Limber approximation, which sets $\chi_{1}=\chi_{2}$, works well especially for higher $\ell$’s, where the feature that the integrands have for $\chi_{1}=\chi_{2}$ gets increasingly sharp. In addition, for high multipoles $\ell$ the flat sky approximation is valid. Moreover, the approximation works best when the kernels $K_{i}^{\mathrm{A}}$ are broad in comoving distance (i.e., the approximation works best for the weak lensing sample rather than the clustering one) and for the auto-correlation within a redshift bin, as the kernels $K_{i}^{\mathrm{A}}$, $K_{j}^{\mathrm{A}}$ are superimposed in comoving distance.
@@
$C^{\mathrm{AB}}_{ij}(\ell)=N(\ell)\int\frac{\mathrm{d}\chi}{\chi^{2}}K_{i}(\chi)K_{j}(\chi)P_{\mathrm{AB}}\left(k=\frac{\nu}{\chi}\right)\Biggl{\{}1\\ +\frac{1}{\nu^{2}}\Biggl{[}\frac{\chi^{2}}{2}\left(\frac{\tilde{K}_{i}^{{}^{\prime\prime}}(\chi)}{\tilde{K}_{i}(\chi)}+\frac{\tilde{K}_{j}^{{}^{\prime\prime}}(\chi)}{\tilde{K}_{j}(\chi)}\right)\\ +\frac{\chi^{3}}{6}\left(\frac{\tilde{K}_{i}^{{}^{\prime\prime\prime}}(\chi)}{\tilde{K}_{i}(\chi)}+\frac{\tilde{K}_{j}^{{}^{\prime\prime\prime}}(\chi)}{\tilde{K}_{j}(\chi)}\right)\Biggr{]}\Biggr{\}}.$ (13)
@@ 3.1 The Algorithm
$C_{ij}^{\mathrm{gg}}(\ell)=\frac{2}{\pi}\int_{0}^{\infty}\mathrm{d}\chi_{1}\,W_{i}^{g}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}\,W_{j}^{g}(\chi_{2})\\ \times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})j_{\ell}(k\chi_{1})j_{\ell}(k\chi_{2}).$ (15)
@@
$C_{ij}^{\mathrm{ss}}(\ell)=\frac{2(\ell+2)!}{\pi(\ell-2)!}\int_{0}^{\infty}\mathrm{d}\chi_{1}\,W_{i}^{s}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}\,W_{j}^{s}(\chi_{2})\\ \times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})\frac{j_{\ell}(k\chi_{1})}{(k\chi_{1})^{2}}\frac{j_{\ell}(k\chi_{2})}{(k\chi_{2})^{2}}.$ (17)
@@
$W_{i}^{s}(\chi)=\frac{3H_{0}^{2}\Omega_{m}}{2a}\chi\int_{z}^{\infty}\mathrm{d}z^{\prime}n_{i}(z^{\prime})\frac{\chi(z^{\prime})-\chi}{\chi(z^{\prime})}.$ (18)
@@
$C_{ij}^{\mathrm{gs}}(\ell)=\frac{2}{\pi}\sqrt{\frac{(\ell+2)!}{(\ell-2)!}}\int_{0}^{\infty}\mathrm{d}\chi_{1}\,W_{i}^{g}(\chi_{1})\int_{0}^{\infty}\mathrm{d}\chi_{2}\,W_{j}^{s}(\chi_{2})\\ \times\int_{0}^{\infty}\mathrm{d}k\,k^{2}P_{\mathrm{AB}}(k,\chi_{1},\chi_{2})\frac{j_{\ell}(k\chi_{1})}{(k\chi_{1})^{2}}j_{\ell}(k\chi_{2}).$ (19)
@@
where $T_{n}(x)$ are the Chebyshev polynomials, and the coefficients $a_{n}$ are given by:
@@
In practice, it is impossible to compute an infinite number of coefficients $a_{n}$ thus, $f(x)$ is approximated by truncating the series to a finite sum:
$\tilde{f}(x)=\sum_{n=0}^{N-1}a_{n}T_{n}(x).$ (24)
Alternatively, $f(x)$ can be approximated using interpolation at Chebyshev points, a set of points defined as:
$t_{n}=\cos{\left(\frac{2n+1}{2N}\pi\right)}\,\,\,n=0,...\,,N-1.$ (25)
Those points are defined in $[-1,1]$ and have the property of being more dense close to the edges of the interval. The interpolating polynomial at $N$ Chebyshev points is given by:
$\hat{f}(x)=\sum_{n=0}^{N-1}c_{n}T_{n}(x).$ (26)
While the two approximations defined in Eq. (24) and (26) yield similar results, they are not identical. One approximation is obtained by the truncation of the infinite series expansion on the basis of the Chebyshev polynomials, the other one by interpolation in Chebyshev points. As a consequence, $a_{n}$ and $c_{n}$ are different sets of coefficients. However, using the properties of Chebyshev polynomials, and the fact that:
$\forall n\in\{0,N-1\}\quad\hat{f}\left(t_{n}\right)=f\left(t_{n}\right),$ (27)
one can show (Trefethen, 2019) that they are related:
$c_{n}=a_{n}$ $+(a_{n+2k}+a_{n+4k}+\dots)$
$+(a_{-n+2k}+a_{-n+4k}+\dots).$ (28)
for $11\leq n\leq k-$ 1. That is, the coefficients in the infinite series expansion are reassigned to their aliases of degree $<k$. This property explains why Chebyshev polynomials are widely used in approximation theory, as they offer greater accuracy than other methods for a fixed number of interpolation points. As shown in Trefethen (2019), Chebyshev polynomials are nearly optimal approximants, with a negligible difference in accuracy compared to interpolation at truly optimal points. Moreover, the slight improvement in accuracy does not justify the additional computational cost of determining and using the optimal interpolant. In this specific problem, we approximate the power spectrum by interpolation in the Chebyshev points (Eq. (26)). This choice is driven by the fact that the coefficients $c_{n}$ can be evaluated through a Discrete Fourier Transform (DFT) (Press, 2007; Frigo & Johnson, 1997):
$c_{n}=\sum_{j=0}^{N-1}f(x_{j})\mathrm{e}^{2\mathrm{i}\pi jn/N}.$ (29)
Computing an FFT is much more efficient than evaluating the integrals defining the coefficients $a_{n}$ (Eq. (23)), which justifies our approach.
The power spectrum, a smooth and continuous function over the interval $[k_{\mathrm{min}},k_{\mathrm{max}}]$ (which can be remapped in $[-1,1]$), in the basis of the Chebyshev polynomials becomes:
$P(k,\chi_{1},\chi_{2};\theta)\approx\sum_{n=0}^{n_{\mathrm{max}}}c_{n}(\chi_{1},\chi_{2};\theta)T_{n}(k),$ (30)
where the dependency on a generic set of cosmological parameters, $\theta$, has now been made explicit. The choice of the sampling points $k_{j}$ for the power spectrum will be motivated in the next section. This decomposition enables a clear separation of geometric and cosmological components in the integrals, simplifying the overall calculation. We can write:
@@
$f^{\mathrm{AB}}(k)=\begin{cases}k^{2}&\mathrm{AB}=\mathrm{gg}\,,\\ 1/k^{2}&\mathrm{AB}=\mathrm{ss}\,,\\ 1&\mathrm{AB}=\mathrm{gs}\,.\end{cases}$ (33)
@@
$C_{ij}^{\mathrm{AB}}(\ell)=\int_{0}^{\infty}\mathrm{d}\chi\int_{0}^{1}\mathrm{d}R\,\chi\Bigl{[}\mathcal{K}_{i}^{\mathrm{A}}(\chi)\mathcal{K}_{j}^{\mathrm{B}}(R\chi)\\ +\mathcal{K}_{j}^{\mathrm{B}}(\chi)\mathcal{K}_{i}^{\mathrm{A}}(R\chi)\Bigr{]}w_{\ell}^{\mathrm{AB}}(\chi,R\chi).$ (34)